Reduce the following equation using Boolean algebra and show all of your steps.

Q = A'B'C'D' + A'B'C'D + A'B'CD' + AB'C'D' + AB'C'D + AB'CD' + AB'CD
Answer #1

ANSWER :--

Giver, boolean as :- equation Q = Aeco+ Aecd + ABCA + ABeb + ABCD A B C D + ABCD & implifying above using usin+ B = ~ = = = B B c blos (0+A)} Bc + Bc ot ocA B (d+CA) + BCD! 3pCc+c) {c+a)} + cor [ucing p (1) (c+a)} + BCB! (ct 4 AB + CD) C+c) (c+6) } AAB = [ using x+y2 = (x+4) (x+2)] B{ ) (c+b)} + AB! + AB [uring x+x= 1] B (c+b) +